OpenStudy (anonymous):

If f(x) = 9 – 4|x + 1|, find f(–5). A. –15 B. –7 C. 7 D. 23 I got A.. Is that right?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

try agian

OpenStudy (anonymous):

Is it C?

ganeshie8 (ganeshie8):

\( f(x) = 9 – 4|x + 1|\) \( f(-5) = 9 – 4|-5+ 1|\) \( ~~~~~~~~~ = 9 – 4|-4|\) \( ~~~~~~~~~ = 9 – 4*4\) \( ~~~~~~~~~ = 9 – 16\) \( ~~~~~~~~~ = -7\)

OpenStudy (anonymous):

Ohh ok! I forgot to put the negative on it!

OpenStudy (anonymous):

Thanks!

ganeshie8 (ganeshie8):

I see, u wlc :)